OpenStudy (anonymous):

What is the x−intercept and y−intercept of the graph of 3x − 4y = 12? x−intercept = (−4, 0); y−intercept = (0, 3) x−intercept = (4, 0); y−intercept = (0, −3) x−intercept = (0, −3); y−intercept = (4, 0) x−intercept = (0, 3); y−intercept = (−4, 0)

jimthompson5910 (jim_thompson5910):

Hint: To find the x-intercept, plug in y = 0 and solve for x To find the y-intercept, plug in x = 0 and solve for y

jimthompson5910 (jim_thompson5910):

I'll do the x-intercept to get you started 3x − 4y = 12 3x − 4(0) = 12 ... plug in y = 0 3x - 0 = 12 3x = 12 x = 12/3 x = 4 So the x-intercept is (4,0)